Using de Moivres theorem, express

Cos5A and sin4A as polynomials in sin A and cos A.
1
Expert's answer
2020-05-04T06:27:09-0400

De Moivre's Theorem:

For any complex number "z=r(\cos \theta +i \sin \theta)" and any integer "n",

"z^n=r^n(\cos (n\theta)+i \sin(n\theta))"


We will consider "z=\cos A+i \sin A"

"z^5=\cos 5A+i \sin 5A"

"z^5=(\cos A+i \sin A)^5=\cos^5A+5\cos ^4 A(i \sin A)+10\cos ^3 A(i \sin A)^2+10\cos ^2A(i \sin A)^3+5\cos A(i\sin A)^4+(i \sin A)^5=(\cos ^5A-10\cos ^3 A\sin^2 A+5\cos A\sin^4A)+i(5\cos^4 A\sin A-10\cos^2 A\sin ^3 A+\sin^5A)"

Therefore, "\cos 5A= \cos ^5A-10\cos ^3 A\sin^2 A+5\cos A\sin^4A"


"z^4=\cos 4A+i \sin 4A"

"z^4=(\cos A+i \sin A)^4=\cos ^4A+4\cos ^3A(i\sin A)+6\cos^2 A(i\sin A)^2+4\cos A(i \sin A)^3+(i\sin A)^4=(\cos^4A-6\cos^2 A\sin^2A+\sin^4A)+i(4\cos ^3A\sin A-4\cos A\sin^3A)"

Therefore, "\sin 4A= 4\cos ^3A\sin A-4\cos A\sin^3A"


Answer: "\cos 5A= \cos ^5A-10\cos ^3 A\sin^2 A+5\cos A\sin^4A"

"\sin 4A= 4\cos ^3A\sin A-4\cos A\sin^3A"